OpenStudy (alivejeremy):

Chris’s hike started at an elevation of –12 m. His hike finished at an elevation of 235 m. How much did Chris’s elevation change?

OpenStudy (alivejeremy):

can someone help plz

OpenStudy (alivejeremy):

u get it

OpenStudy (sagewilson):

is this an addition question? If so the answer would be "Chris's elevation changed by 247m"

OpenStudy (alivejeremy):

yes ur right

OpenStudy (alivejeremy):

can u help wit more if u can and i'll help u if i can

OpenStudy (sagewilson):

sure.

OpenStudy (alivejeremy):

ok

OpenStudy (alivejeremy):

u got any qusetion

OpenStudy (alivejeremy):

questions

OpenStudy (alivejeremy):

Ravi had $11 in his piggy bank on Monday. He deposited $1.30 into the piggy bank on Tuesday and removed $2.85 on Wednesday. How much was left in Ravi’s piggy bank?

OpenStudy (alivejeremy):

u still there

OpenStudy (sagewilson):

yeah

OpenStudy (alivejeremy):

ok lol

OpenStudy (alivejeremy):

i think its 9.45

OpenStudy (sagewilson):

you're correct.

OpenStudy (alivejeremy):

Nolan lost 12.5 points in the final round of a game, bringing his score down to –8. Suppose p represents Nolan’s points before he lost the final round. Which equation can be used to represent this situation? A. p – (–8) = –12.5 B. p + 12.5 = –8 C. p – 12.5 = –8 D. p + 8 = 12.5

OpenStudy (alivejeremy):

and i don't have one idea

OpenStudy (sagewilson):

it's okay. Just give me a minute.

OpenStudy (alivejeremy):

i think its C

OpenStudy (sagewilson):

Yup. It is C.

OpenStudy (alivejeremy):

On Wednesday, William wrote a check in the amount of $45.60. This brought his balance down to –$15.75. Suppose n represents William’s balance before he wrote the check. Which equation can be used to represent this situation? A. n + 15.75 = –45.60 B. n – 15.75 = 45.60 C. n + 45.60 = –15.75 D. n – 45.60 = –15.75

OpenStudy (sagewilson):

It is D.

OpenStudy (alivejeremy):

yup its D

OpenStudy (alivejeremy):

ok u got any

OpenStudy (alivejeremy):

i can help u

OpenStudy (sagewilson):

Not at the moment.

OpenStudy (alivejeremy):

lol

OpenStudy (sagewilson):

But if you need anymore help just message me.

OpenStudy (alivejeremy):

ok kool

OpenStudy (alivejeremy):

i just fan u

OpenStudy (sagewilson):

Cool. You might wanna close this up so you can ask questions later.

OpenStudy (alivejeremy):

ok

OpenStudy (alivejeremy):

ok its closed